A large food-processing corporation is considering using laser technology to speed up and eliminate waste in the potato-peeling process. To implement the system, the company anticipates needing RM4 million to purchase the industrial-strength lasers. The system will save income taxes will also increase by RM160,000. The system is expected to have a 9-ycar service life and will have a salvage value of about RM200,000. If the company s MARR s l 6%, use the NPW method tojustify the pro ect. PW of the system Is the system justified to be implemented? Format: -224005 Format: xsupposed to be same format
Answer #1

Annual net savings = Annual revenue - Annual cost - Annual tax = 1,350,000 - 360,000 - 160,000 = 830,000

(a)

NPW = - 4,000,000 + 840,000 x P/A(16%, 9) + 200,000 x P/F(16%, 9)

= - 4,000,000 + 840,000 x 4.6065 + 200,000 x 0.2630

= - 4,000,000 + 3,869,460 + 52,600

= - 77940

(b) NO.

Since NPW < 0, system is not justified.